OpenStudy (anonymous):

The first peak of a roller coaster is 55m above the ground. The 1200 kg car starts from rest and goes down the hill and up the second hill which is 30m high. How fast is the car traveling at the top of the second hill?

OpenStudy (shamim):

wt do u wanna know frm this math is not clear to me

OpenStudy (shamim):

now i m trying to solve ur problem in my own way

OpenStudy (shamim):

the final velocity of the roller coaster after coming frm 1st pick is

OpenStudy (shamim):

\[v ^{2}=u ^{2}+2gh\]

OpenStudy (shamim):

here u=0, g=10m/s, h=55m

OpenStudy (shamim):

\[or, v ^{2}=0^{2}+2 \times 10 \times 55\]

OpenStudy (anonymous):

2,200 m/s is the answer?

OpenStudy (shamim):

\[or, v=\sqrt{2 \times 10 \times 55}=33ms ^{-1}\]

OpenStudy (anonymous):

so the answer is 2,200 m/s

OpenStudy (shamim):

still not

OpenStudy (anonymous):

now then this question uses your good old. KE+PE = KE+PE. since there is no work done from external force, the change in KE = change in PE \(\frac{1}{2}mv^2=mg(h-h_o)\) rearranging, \(v=\sqrt{2g(h-h_o)}\) i.e. not 2200 do you need further help?

OpenStudy (shamim):

we hv to find out the velocity of the roller coaster when it is on the 2nd pick

OpenStudy (anonymous):

yes is \[v=\sqrt{2g(h-h _{o})}\] the answer to the problem?

OpenStudy (anonymous):

ah ok. soo im lost xD

OpenStudy (shamim):

@shadow............. ur system is easier than mine

OpenStudy (anonymous):

yup, derived from your original equation. mechanical energy is conserved. \(KE_i + PE_i =KE_f +PE_f\) rearranging, \(PE_f - PE_i =KE_f -KE_i\) since \(KE_i\) =0, (the body starts from rest) \(mgh-mgh_o=\frac{1}{2} mv^2\) rearranging again, \(v=\sqrt{2g(h-h_o)}\)

OpenStudy (anonymous):

ok so we are trying to find \[v _{f}\] rite?

OpenStudy (anonymous):

yup! :)

OpenStudy (anonymous):

Note:The entire journey has so other force acting on it except a conservative force, gravity. So, Mechanical energy is conserved.

OpenStudy (anonymous):

so how do we find \[v _{f}\]

OpenStudy (anonymous):

use your equation for kinetic energy and potential energy. the one you normally post after your questions.

OpenStudy (anonymous):

im still lost..

OpenStudy (shamim):

here h=55m

OpenStudy (shamim):

\[h _{0}=30m\]

OpenStudy (anonymous):

so i have to find the first velocity first before ifind the other velocity on the 2nd peak?

OpenStudy (shamim):

g=9.8m/s

OpenStudy (anonymous):

yes. Now, the first velocity is zero, though. so, initial KE=0

OpenStudy (anonymous):

so the equation Shadowys gave me the v=22.14m/s?

OpenStudy (shamim):

i think the result will b 2214cm/s

OpenStudy (anonymous):

Yup. though did you understand why it is so?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oh so the answer to this problem is 2214cm/s because shamim converted meters to cm? xD

OpenStudy (shamim):

ya i converted it frm m to cm

OpenStudy (anonymous):

so the speed in which the car is traveling at the top of the second hill is 22.14m/s or 2214cm/s?

OpenStudy (anonymous):

lol I guess you knew. I believe you understood why v is calculated that way?

OpenStudy (shamim):

anyway i can solve this problem in my way too as i started b4

OpenStudy (anonymous):

so the answer to this problem is 22.14m/s or 2214cm/s?

OpenStudy (shamim):

ya

OpenStudy (anonymous):

ok ty :)

OpenStudy (shamim):

ok i m going to start

OpenStudy (shamim):

1st i wanna tell abt 1st pick

OpenStudy (anonymous):

yup. There will be a complication if we simply use the motion formula \(v^2 =u^2 +2aS\) though, since this question require a scalar answer, your answer might be a little off since it's a vector eq instead...you can try though.

OpenStudy (shamim):

u know the velocity of roller coaster on 1st pick is 0

OpenStudy (anonymous):

sii C:

OpenStudy (anonymous):

that means yes xD

OpenStudy (anonymous):

lol @lilfayfay you french?

OpenStudy (anonymous):

no, i take Spanish thats why :D

OpenStudy (shamim):

we can use this equation for scalar too

OpenStudy (anonymous):

okie dokie :)))

OpenStudy (anonymous):

oh right. I just checked. same answer. lol I guess i'm too careful.

OpenStudy (shamim):

so let the final velocity of roller coaster after coming from 1st pick to the ground=v

OpenStudy (shamim):

\[v ^{2}=u ^{2}+2gh\]

OpenStudy (shamim):

since u=0

OpenStudy (shamim):

did u understand why u=0?

OpenStudy (anonymous):

because velocity is 0

OpenStudy (shamim):

\[or, v ^{2}=0^{2}+2 \times 10 \times 55\]

OpenStudy (anonymous):

so why is u=0?

OpenStudy (shamim):

\[or, v=\sqrt{2 \times 10 \times 55}=33m\]

OpenStudy (shamim):

u=o, because the roller coaster was at rest when it was on 1st pick

OpenStudy (shamim):

rest means velocity is 0

OpenStudy (anonymous):

ok i thought v=22.14m/s

OpenStudy (shamim):

not now

OpenStudy (shamim):

we hv to do more math to get result

OpenStudy (shamim):

now i wanna do math for pick 2

OpenStudy (shamim):

initial velocity u=33m/s

OpenStudy (anonymous):

but Shadowys did an easier way?

OpenStudy (shamim):

final velocity of roller coaster when it is on the 2nd pick=v

OpenStudy (shamim):

ya his way is easier than mine

OpenStudy (anonymous):

oh, ok :)

OpenStudy (shamim):

i think u should know all the ways to solve ur math. every way has some understanding

OpenStudy (shamim):

knowledge

OpenStudy (shamim):

should i proceed or not

OpenStudy (anonymous):

proceed but im going to bed in a few mins since its late here lol

OpenStudy (shamim):

lol

OpenStudy (shamim):

i think u understood my way